OpenStudy (cp9454):

How many kg of sodium and liquid chlorine can be obtained from 585 metric ton of salt ?

OpenStudy (vincent-lyon.fr):

Sodium: 21 g/mol Chlorine: 35.5 g/mol Work out mass ratios.

OpenStudy (abmon98):

Molar mass Na = 23.0g/mol Molar mass Cl = 35.5g/mol Molar mass NaCl = 58.5g/mol NaCl → Na+ + Cl- 1mol NaCl will produce 1 mol Na and 1 mol Cl Mass of Na produced from 585metric tonnes NaCl = 23/85.5*585 = 230 metric tonnes of sodium *1000 to change into Kg Mass of Cl produced from 585 metric tonnes NaCl = 35.5/85.5*585 = 355 metric tonnes chlorine *1000 to change to Kg

OpenStudy (vincent-lyon.fr):

Sorry for my typo. Sodium is 23 g/mol as Abmon wrote.

OpenStudy (jfraser):

dont forget that elemental chlorine is \(diatomic\), so the balanced reaction, which everyone has skipped, is \[2NaCl \rightarrow 2Na + Cl_2\]